40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Nov 1 16:06:32 2014, Koji, Configuration, LSC, DRMI locked 6x
    Reply  Sat Nov 8 16:01:32 2014, ericq, Configuration, LSC, DRMI sensing 6x
       Reply  Mon Nov 10 18:23:10 2014, ericq, Configuration, LSC, DRMI sensing REFL11.pngREFL33.pngREFL55.pngREFL165.pngAS55.png
Message ID: 10661     Entry time: Sat Nov 1 16:06:32 2014     Reply to this: 10690
Author: Koji 
Type: Configuration 
Category: LSC 
Subject: DRMI locked 

Continued from ELOG 10659


DRMI locking

Following Jenne's elog entry in Aug 2013 (9049), DRMI was configured and locked. The lock was stable, indefinite, and repeatitive.

- DRMI Configuration

Demod phases has not been changed from PRMI

REFL11: WTN 0dB PHASE 21deg, REFL11I x0.1 -> PRCL
REFL55: WTN 21dB PHASE 25deg, REFL55Q x1 -> MICH, REFL55I x1 -> SRCL

AS110 phase was adjusted to maximize Q during the lock: +1deg (AS110Q_ERR was +4400 ~ +5500)

PRCL: GAIN -0.05 FM4/5 ON, Triggered FM 2/3/6/9, Servo trigger: POP22I 20up 10down, No Normaization.
MICH: GAIN +1 FM4/5 ON, Triggered FM 2/3/6/9, Servo trigger: POP22I 20up 10down, No Normaization.

SRCL: GAIN +2 FM4/5 ON, Triggered FM2/3/6/8/9, Servo trigger: AS110Q up 500 down 5, No Normaization.
(FM8 was set to be x2.5 flat gain such that the gain is increased after the lock)

MICH actuation is still BS+PRM and does not include SRCL decoupling yet.
This should be fixed ASAP.

DRMI Calibration

Let's use these entries 

SRM: http://nodus.ligo.caltech.edu:8080/40m/10664
SRM = (19.0 +/- 0.7) x 10 -9/ f2

PRM: http://nodus.ligo.caltech.edu:8080/40m/8255
PRM:  (19.6 +/- 0.3) x 10 -9 / f2 m/counts

BS/ITMs http://nodus.ligo.caltech.edu:8080/40m/8242
BS     = (20.7 +/- 0.1)    x 10 -9 / f2 m/counts
ITMX = (4.70 +/- 0.02)  x 10 -9/ f2
m/counts
ITMY = (4.66 +/- 0.02) x 10 -9/ f2
m/counts

- PRCL Calibration

Lock-in oscillator module 675.13Hz 100 -> +1 PRM

Measurement bandwidth 0.1Hz -> Signal power BW 0.471232 (FLATTOP window)

C1:SUS-PRM_LSC_IN1: 97.45 cnt/rtHz => 4.19 pm/rtHz

REFL11I: 12.55   cnt/rtHz => 3.00e12 cnt/m
REFL11Q:  0.197  cnt/rtHz => 4.70e10 cnt/m
=> 0.90 deg rotated! (GOOD)

REFL33I:  1.63   cnt/rtHz => 3.89e11 cnt/m
REFL33Q:  0.196  cnt/rtHz => 4.68e10 cnt/m
=> 8.32 deg rotated!

REFL55I:  0.0495 cnt/rtHz => 1.18e10 cnt/m
REFL55Q:  0.548  cnt/rtHz => 1.31e11 cnt/m => 84.8 deg rotated! (WHAT!)

REFL165I: 1.20   cnt/rtHz => 2.86e11 cnt/m
REFL165Q: 0.458  cnt/rtHz => 1.09e11 cnt/m
=> 20.9 deg rotated!

- MICH Calibration

Lock-in oscillator module 675.13Hz 100 -> -1 ITMX +1 ITMY

Measurement bandwidth 0.1Hz -> Signal power BW 0.471232 (FLATTOP window)

C1:SUS-ITMX_LSC_IN1: 121.79 cnt/rtHz => 1.26pm/rtHz
C1:SUS-ITMY_LSC_IN1: 121.79 cnt/rtHz => 1.25pm/rtHz

AS55Q:   12.45   cnt/rtHz => 4.96e12 cnt/m (STRONG)

REFL11I:  0.0703 cnt/rtHz => 2.80e10 cnt/m
REFL11Q:  0.0142 cnt/rtHz => 5.66e09 cnt/m
=> 78.5 deg rotated! (WHAT!)

REFL33I:  0.0473 cnt/rtHz => 1.88e10 cnt/m
REFL33Q:  0.0291 cnt/rtHz => 1.16e10 cnt/m => 58.4 deg rotated!

REFL55I:  0.00668cnt/rtHz => 2.66e09 cnt/m
REFL55Q:  0.0261 cnt/rtHz => 1.04e10 cnt/m => 14.4 deg rotated! (OK)

REFL165I: 0.0233 cnt/rtHz => 9.28e09 cnt/m
REFL165Q: 0.0512 cnt/rtHz => 2.04e10 cnt/m => 24.5 deg rotated! (GOOD)

- SRCL Calibration

Lock-in oscillator module 675.13Hz 100 -> SRM

Measurement bandwidth 0.1Hz -> Signal power BW 0.471232 (FLATTOP window)

C1:SUS-SRM_LSC_IN1: 121.77 cnt/rtHz => 5.08pm/rtHz

AS55I:    0.256   cnt/rtHz => 5.05e10 cnt/m
AS55Q:    0.3498  cnt/rtHz => 6.90e10 cnt/m

REFL11I:  0.00624 cnt/rtHz => 1.23e09 cnt/m
REFL11Q:  0.00204 cnt/rtHz => 4.02e08 cnt/m

REFL33I:  0.00835 cnt/rtHz => 1.65e09 cnt/m
REFL33Q:  0.0659  cnt/rtHz => 1.30e10 cnt/m

REFL55I:  0.0201  cnt/rtHz => 3.97e09 cnt/m
REFL55Q:  0.01505 cnt/rtHz => 2.97e09 cnt/m

REFL165I: 0.0238  cnt/rtHz => 4.69e09 cnt/m
REFL165Q: 0.0247  cnt/rtHz => 4.87e09 cnt/m

DRMI Openloop measurements
Servo filter TF measurements

The UGFs were ~250Hz for PRCL and ~100Hz for MICH, and ~250Hz for SRCL, respectively.
MICH showed (presumably) crosscoupling related peak ~350Hz. SRCL had small deviation from the model.
This may also be related to the cross couplig.

The OLTF was modelled by the servo and violin filters TF from foton, estimated TF of the AA/AI filters, and the constant time delay.

Displacement spectra measurement

- PRCL

The OLTF compensation was not actually succesfull at 300Hz, but otherwise the situation is very similar to the one with PRMI.

- MICH

Again the servo compensation at 300Hz was not successful. If we believe that AS55Q is the best MICH sensor, the out-of-loop
noise level of MICH was quite similar to the one in PRMI. We should try to use AS55Q for DRMI MICH for investigation purpose
to see which REFL signal has the best MICH quality. REFL165 seems to be iproved in the signal amplitude. Can we use this
for locking now?

- SRCL

It is in fact difficult to tell what is the correct out-of-loop noise level. AS55I has too much contamination from MICH and is not indicating
useful info. This measurement should be tried once the sensor diagonalization is done.

REFL55I is not seeing anything real abobe 30Hz. We should be able to reduce the UGF and the servo gain.

The absolute motion level of SRCL is something similar to PRCL, rather than MICH.

 

Attachment 1: DRMI_PRCL_OLTF.pdf  1.364 MB  Uploaded Mon Nov 3 19:18:42 2014  | Hide | Hide all | Show all
DRMI_PRCL_OLTF.pdf
Attachment 2: DRMI_MICH_OLTF.pdf  1.401 MB  Uploaded Mon Nov 3 19:19:13 2014  | Show | Hide all | Show all
Attachment 3: DRMI_SRCL_OLTF.pdf  1.370 MB  Uploaded Mon Nov 3 19:22:09 2014  | Hide | Hide all | Show all
DRMI_SRCL_OLTF.pdf
Attachment 4: DRMI_PRCL_SPE.pdf  888 kB  Uploaded Mon Nov 3 20:34:19 2014  | Hide | Hide all | Show all
DRMI_PRCL_SPE.pdf
Attachment 5: DRMI_MICH_SPE.pdf  2.035 MB  Uploaded Mon Nov 3 20:36:00 2014  | Show | Hide all | Show all
Attachment 6: DRMI_SRCL_SPE.pdf  1.971 MB  Uploaded Mon Nov 3 20:36:50 2014  | Hide | Hide all | Show all
DRMI_SRCL_SPE.pdf
ELOG V3.1.3-